7. _Pb + K2SO4 → PbSO4 + K How many atoms of potassium are produced from 2.50 * 10” units of K2SO4
Answer #1

Answer -

Given,

K2SO4 = 2.50 * 1023 units

Number of atoms produced = ?

_____ Pb + ______ K2SO4\rightarrow ______ PbSO4 + _____ K

First Balance the reaction,

For Balancing a reaction make the number of atoms of each species equal on both side.

1 Pb + 1 K2SO4\rightarrow 1 PbSO4 + 2 K [BALANCED]

Now,

According to Avogadro's Law,

1 mole of anything have 6.022 * 1023 atoms, molecules or ions

i.e. 1 atom, molecule or ion is equal to (1/6.022 * 1023) mol

So,

2.50 * 1023 units of K2SO4 is equal to, 2.50 * 1023 * (1/6.022 * 1023) mol

So, moles of K2SO4 = 0.41514447027 mol

1 Pb + 1 K2SO4\rightarrow 1 PbSO4 + 2 K [BALANCED]

Using Stiochiometry,

It can be analyzed that for 1 mol of K2SO4, 2 moles of K are produced.

i.e. moles of K Produced = 2 * moles of K2SO4

So, Moles of K produced = 2 * 0.41514447027 mol = 0.83028894054 mol

Using Avogadro's Law,

1 mole = 6.022 * 1023 atoms

So, 0.83028894054 mol of K = 0.83028894054 * 6.022 * 1023 atoms = 4.99999999993 * 1023 atoms

So, Approx. 5 * 1023 atoms of K are produced [ANSWER]